OpenStudy (anonymous):

Let a_n=2_n+6 and let S_n=a_1+a_2+a_3…+a_n find a_(n ) and s_n when n=30

OpenStudy (anonymous):

Wouldn't a of n =2(30)+6

OpenStudy (anonymous):

No that would be only the 30th term of the sequence.

OpenStudy (anonymous):

and that is 66 right

OpenStudy (anonymous):

so what is Sn?

OpenStudy (anonymous):

\[S _{30}= \sum_{1}^{30}2n+6\]

OpenStudy (anonymous):

Oh and I didn't see that they asked for a30 yeah that's the correct way to find it.

OpenStudy (anonymous):

I'll help you calculate S30 in just a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[S _{30}=\sum_{1}^{30}2n+6=2\sum_{1}^{30}n+\sum_{1}^{30}6\] now you need to know that \[\sum_{1}^{k}n=k(k+1)/2\]and\[\sum_{1}^{k}c=kc\]Can you take it from here?

OpenStudy (anonymous):

too complex

OpenStudy (anonymous):

What particularly do you not understand?

OpenStudy (anonymous):

kindly solve for Sn

OpenStudy (anonymous):

Do you understand what I did to break up the 2n+6 to two different sums?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Ok that just comes from the linearity of summations. Basically sum(x+y)=sumx+sumy and sum(cx)=csum(x).

OpenStudy (anonymous):

I should have mentioned in that property earlier that the c was a constant.

OpenStudy (anonymous):

Hmm, take Sn=30/2(2a+(30-1)d) Can it be solved this way and if then what is d?

OpenStudy (anonymous):

what is a and d ?

OpenStudy (anonymous):

What is that equation?

OpenStudy (anonymous):

formula for finding the sum of a1+a2+a3... a of n

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/partial-sums.html Check out this site. It is where I am getting these identities from.

OpenStudy (anonymous):

Sn=n/2(2a+(n-1)d)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'll write out the steps of the solution step by step while you read. I'll try to show as much work as possible so feel free to ask questions!

OpenStudy (anonymous):

\[S _{30}=\sum_{1}^{30}2n+6=\sum_{1}^{30}2n+\sum_{1}^{30}6\] Then we can pull the 2 out of the first term:\[=2\sum_{1}^{30}n+\sum_{1}^{30}6\]Now I will evaluate the two sums based on those identities I posted earlier.\[=2(30)(30+1)/2+6(30)=930+180=1110\]